OpenStudy (anonymous):

For each equation, find the equation that is (a) parallel and (b) perpendicular that passes through the specified point. 1. y = x - 1

OpenStudy (anonymous):

Can you help!!?

OpenStudy (anonymous):

parallel: y=x+k perpendicular: y=-x+k

OpenStudy (anonymous):

here, k is any constant value

OpenStudy (anonymous):

What I got so far is y = x - 1 with plot point of (2,0) the formula my instructor gave me is Y = mx + b working from that I get 0 = -1(2) = b then 0 = -2 = b and then 2 = b which is the perpendicular line but I am having trouble fining the parallel line If it is the negative recip would it be -2 = b for the parraellel then?